OpenStudy (lizabethliong):

The sum of two number is 55. The quotient obtained by dividing the larger number by the smaller number is 2, with a remainder of 7. Find the difference in the reciprocals of the numbers. The final answer has to be 23/624 Thanks in advance!

OpenStudy (jiteshmeghwal9):

Let the larger number be 'a' & smaller number be 'b'. Then According to question, a=2b+7 (using Euclid's division algorithm ) a+b=55

OpenStudy (jiteshmeghwal9):

solve for a & b

OpenStudy (jiteshmeghwal9):

one way to solve this type of questions is to do as the question is asking to u step by step

OpenStudy (lizabethliong):

a=39 b=16 the next step please?

OpenStudy (phi):

Find the difference in the reciprocals of the numbers. the reciprocal of x is 1/x (i.e. you "flip" the number) in this problem, find \[ \frac{1}{16}- \frac{1}{39} \]